OpenStudy (anonymous):

Find the product of (x + 3)^2

OpenStudy (aliahmed):

x^2-6x+9

jhonyy9 (jhonyy9):

aliahmed can you prove it easy but i think that this not is correct,right check it please

jhonyy9 (jhonyy9):

(a+b)(a+b)=a^2 +ab +ab +b^2 =a^2 +2ab +b^2 ok ?

OpenStudy (aliahmed):

yes you are right iam sorry

OpenStudy (anonymous):

(x + 3)^2 = x^2 +3x +3x +9 = x^2 +6x +9

jhonyy9 (jhonyy9):

attention please good luck bye

OpenStudy (aliahmed):

thankx bye